OpenStudy (infinexa):

Rebecca had 9/10 of a yard of ribbon. Then she gave 7/20 of a yard to a friend. How much does she have left?

OpenStudy (infinexa):

@45h13yshorty

OpenStudy (infinexa):

@Rushwr

OpenStudy (infinexa):

@BigJJRob

OpenStudy (rushwr):

\[\frac{ 9 }{ 10 } - \frac{ 7 }{ 20}\] First we have to bring the base numbers to an equal value \[\frac{ 9*2 }{ 10*2 } - \frac{ 7*1 }{20*1 }= \frac{ 18 }{ 20 }- \frac{ 7 }{ 20 }\] when u simplify the equation u will get 11/20 that is the remaining area

OpenStudy (infinexa):

Thank you

OpenStudy (rushwr):

no problem :)